The average kinetic energy of the molecules in a gas sample depends only on the temperature, ?.T. However, given the same kinetic energies, a lighter molecule will move faster than a heavier molecule, as shown in the equation for rms speed

rms speed=3??ℳ⎯⎯⎯⎯⎯⎯⎯⎯⎯√rms speed=3RTℳ

where ?=8.314 J/(mol⋅K)R=8.314 J/(mol⋅K) and ℳℳ is molar mass in kilograms per mole. Note that a joule is the same as a kilogram‑meter squared per second squared (kg·m2/s2).

What is the rms speed of F2F2 molecules at 349 K?
Answer #1

Molar mass of F2,

mm = 38.0g/mol

= 0.038 Kg/mol

T = 349 K

vrms = sqrt (3*R*T/M)

here : T is temperature in Kelvin

and M is molar mass in Kg/mol

So,

vrms = sqrt(3*8.314*349.0/0.038)

= 4.786*10^2 m/s

Answer: 479 m/s